# MeanVC: Lightweight and Streaming Zero-Shot Voice Conversion via Mean Flows

### v1

@@ 1 Introduction
Zero-shot voice conversion (VC) aims to transfer the timbre from the source speaker to any unseen speaker while keeping the linguistic content unchanged. This capability has profound implications for a wide array of applications, including movie dubbing [1, 2], privacy protection [3], personalized pronunciation correction, and the creation of expressive virtual avatars. Benefiting from modern deep learning based approaches, recent voice conversion models [4, 5, 6, 7, 8, 9] have achieved remarkable performance in terms of naturalness and speaker similarity. However, as the demand for real-time processing in various application scenarios increases, the computational cost of high-performing models becomes a significant bottleneck for practical deployment. This has created a pressing need for streaming models that are simultaneously fast (low-latency), lightweight (low computational overhead), and high-fidelity (high audio quality and speaker similarity).
@@
Recent streaming zero-shot VC studies have explored two main architectural paradigms: autoregressive (AR) and non-autoregressive (NAR) frameworks. A representative AR-based streaming VC system is StreamVoice [10], which employs a fully causal, context-aware language model. It alternates between semantic bottleneck features and acoustic codec tokens at each AR step, supported by an auxiliary acoustic predictor, to enable streaming zero-shot VC without future look-ahead. Building on this design, StreamVoice+ [11] introduces a learnable semantic encoder and a connector, while adapting the backbone via LoRA [12]. It further improves quality through a residual-bottleneck connector, forming an end-to-end streaming pipeline. Despite their high-fidelity output, these models fundamentally compromise on the other two critical requirements. Their sequential decoding process inherently leads to high latency, failing the ’fast’ criterion. Moreover, their massive model sizes (101M for StreamVoice and 153M for StreamVoice+) make them far from ’lightweight,’ resulting in prohibitive computational costs that render real-time inference on standard CPUs impractical.
@@
To mitigate the inherent latency of AR models, NAR frameworks have been explored for their potential in parallel generation. DualVC2 [13], for instance, presents a lightweight NAR model based on the conformer architecture, which accelerates inference by incorporating dynamic chunk training. While it achieves faster speeds, its reliance on conventional architectures limits its generalization capability to unseen speakers in zero-shot scenarios. Another NAR approach adapts powerful diffusion models for zero-shot VC. Seed-VC [14] employs a diffusion transformer (DiT) [15] architecture within a sliding-window pipeline. Although this mechanism enables streaming VC, the fixed-size windows can disrupt long-term contextual dependencies and speaker characteristics. Thus, while NAR frameworks successfully address the ’fast’ requirement, they often do so by sacrificing ’high-fidelity,’ either through poor generalization in simpler models or context fragmentation in more complex ones. This leaves a clear gap in the field: a framework that can unite the speed of NAR models with the quality and consistency of their AR counterparts.
@@
To address these challenges, we propose MeanVC, a lightweight framework for streaming zero-shot VC that leverages the advantages of both AR and NAR paradigms while addressing their respective limitations. Specifically, MeanVC employs a chunk-wise autoregressive denoising strategy, which processes audio in individual chunks while explicitly conditioning each chunk on the previous one, thereby preserving long-term speaker consistency. To overcome the latency of iterative diffusion samplers, we introduce mean flows [16], a powerful method that enables the diffusion model to generate high-quality results with 1-NFE (number of function evaluations). We integrate mean flows into the DiT-based decoder. By training the model to regress the average velocity field of the probability flow, MeanVC can synthesize a high-quality spectrogram using 1-NFE. Finally, we incorporate a diffusion adversarial post-training (DAPT) stage to mitigate over-smoothing and further enhance speech quality. With merely 14M parameters, MeanVC achieves superior audio quality and efficiency. Both subjective and objective evaluations demonstrate that MeanVC significantly outperforms existing streaming VC systems. Audio demos and code are publicly available at https://aslp-lab.github.io/MeanVC.
@@ 2.1 Overview
As illustrated in Fig. 1, MeanVC is built on the widely used recognition-synthesis framework [17], consisting of a streaming automatic speech recognition (ASR) module, a speaker encoder, a timbre encoder, a DiT decoder, and a vocoder. First, the pre-trained streaming ASR model extracts bottleneck features (BNFs) from the source waveform. Subsequently, these BNFs are fed into the timbre encoder, which fuses them with fine-grained timbre information extracted from a reference mel-spectrogram to produce timbre BNFs. Meanwhile, speaker embeddings are extracted from a reference waveform via a pre-trained speaker encoder. The DiT decoder generates the converted mel-spectrogram conditioned on both the speaker embeddings and the timbre BNFs. During generation, the source mel-spectrogram is concatenated at the front to serve as a cache, enabling chunk-wise autoregressive denoising and producing the converted spectrogram with the target speaker’s timbre.
@@ 2.2 Chunk-wise Autoregressive Denoising
To enable streaming conversion, we adopt a chunk-wise autoregressive approach that divides the BNFs into smaller chunks and applies a chunk-wise causal mask inspired by MoonCast [18]. This mask strategy allows each chunk to access the historical context from previously generated chunks, ensuring consistency across chunks. Furthermore, we build on conditional flow matching (CFM) to learn optimal transport paths between the two distributions. Given data $x\sim p_{\text{data}}(x)$, prior $\epsilon\sim p_{\text{prior}}(\epsilon)$, the optimal transport flow path can be constructed as: $z_{t}=(1-t)x+t\epsilon$, and the conditional velocity is thus given by $v_{t}=dz_{t}/dt=\epsilon-x$. During training, the objective is to learn a neural network $f_{\theta}$ that minimizes the following CFM loss: $\mathcal{L}_{\text{CFM}}(\theta)=\mathbb{E}_{t,x,\epsilon}\left\|f_{\theta}\left(t,z_{t}\right)-v_{t}\right\|^{2}$.
@@ 2.3 Mean Flows for Generative Modeling
CFM typically requires multiple sampling steps to solve the ordinary differential equation (ODE), and the number of function evaluations directly impacts inference efficiency. With just 1-NFE, the quality of generated spectrograms degrades catastrophically. To address this limitation, we adopt the approach proposed in mean flows [16], which regresses the average velocity field during training to enable high-quality synthesis with only 1-NFE. Specifically, given a time interval $[r,t]$, the displacement on the ODE trajectory within it is $\int_{r}^{t}v(z_{\tau},\tau)d\tau$. The average velocity is then defined as: $u(z_{t},r,t)\triangleq\frac{1}{t-r}\int_{r}^{t}v(z_{\tau},\tau)d\tau$. By differentiating both sides with respect to $t$ and subsequently rearranging the resulting terms, the relationship between $u$ and $t$ can be derived, yielding the mean flows identity:
@@
By utilizing the Jacobian-vector product (JVP), the total derivative can be expressed as: $\frac{d}{dt}u(z_{t},r,t)=v(z_{t},t)\partial_{z}u+\partial_{t}u$. The velocity $v(z_{t},t)$ is the marginal velocity in flow matching. We follow [19] to replace it with the conditional velocity $v_{t}=\epsilon-x$. With this, the target field for training is defined as $u_{\text{tgt}}=v_{t}-(t-r)(v_{t}\partial_{z}u_{\theta}+\partial_{t}u_{\theta})$. We then encourage $f_{\theta}$ to satisfy this field by minimizing the mean flows objective:
@@ 2.4 Diffusion Adversarial Post-Training
Mel-spectrograms generated by flow matching models often suffer from over-smoothing artifacts. While conventional multi-length discriminators [20] are employed, we find that high-frequency components in generated mel-spectrograms still exhibit artifacts such as bright lines. Inspired by [21], we leverage the generator architecture as the discriminator backbone to address this issue. Specifically, we employ the DiT model as generator $G$, which generates spectrograms with 1-NFE. The DiT weights are used to initialize the backbone of the discriminator $D$. After training, the discriminator outputs a logit that distinguishes between real samples $x$ and generated samples $\hat{x}$. To achieve this, we modify the DiT architecture by introducing cross-attention-only Transformer blocks at the second and fourth layers of the 4-layer DiT backbone. These blocks use a learnable vector as the query and perform cross-attention with the backbone’s latent representations as keys and values. The outputs are concatenated channel-wise to form a global feature vector, which is then projected to produce a scalar logit output.
@@ 3.1 Experiment Setup
Dataset. We train MeanVC on the open-source Emilia [22] corpus. Specifically, we use DNSMOS [23] to filter out audio with DNSMOS scores below 3.4, retaining 10,000 hours of mandarin data. All audio files are resampled to 16 kHz for VC training. For zero-shot evaluation, we use the Mandarin subset of the Seed-TTS test set [24], comprising 2,018 source–target testing pairs. For known speakers evaluation, we fine-tune MeanVC on Aishell3 [25], selecting four speakers as targets for evaluation. We use 100 clean and 100 noisy clips as source recordings. To extract semantic features, we incorporate the streaming ASR model Fast-U2++ [26], implemented by WeNet [27] and trained on WenetSpeech [28].
Implementation Details. MeanVC consists of 14M parameters. The DiT decoder employs four DiT blocks, each with a hidden size of 512 and two attention heads. The timbre encoder incorporates two cross-attention modules, with a hidden size of 256 and four attention heads. Fast-U2++ utilizes a 160 ms chunk size for streaming inference and compresses 16 kHz waveforms into semantic features with a 40 ms frame length. Speaker embeddings are extracted using ECAPA-TDNN [29]. We employ Vocos [30] as the vocoder to convert mel-spectrograms to high-fidelity 16 kHz speech waveforms.
@@
Baseline Systems. To validate the performance of MeanVC, we compare MeanVC against the following baseline systems: StreamVoice [10], a language model-based system that employs fully causal context-awareness. It alternately processes semantic and acoustic features to achieve streaming zero-shot voice conversion; Seed-VC [14], a NAR system based on diffusion models. It employs an external timbre shifter to mitigate timbre leakage and achieves streaming zero-shot voice conversion through a sliding-window pipeline; DualVC2 [13], a lightweight NAR system based on the conformer architecture. It uses classic non-causal convolutions and dynamic masking convolutions to achieve streaming any-to-many voice conversion.

### v2

@@ 1 Introduction
Zero-shot voice conversion (VC) aims to transfer the timbre from the source speaker to any unseen speaker while keeping the linguistic content unchanged. This capability has profound implications for a wide array of applications, including movie dubbing [bgmvc, expressive-vc], privacy protection [voiceprivacy2022], personalized pronunciation correction, and the creation of expressive virtual avatars. Benefiting from modern deep learning based approaches, recent voice conversion models [DBLP:conf/icassp/HussainNHLG23, DBLP:conf/icassp/LiLL23, DBLP:conf/icassp/SEF-VC, DBLP:conf/icassp/LuoD24, DBLP:conf/icassp/KimKCNMC25, DBLP:conf/icassp/ChoiP25] have achieved remarkable performance in terms of naturalness and speaker similarity. However, as the demand for real-time processing in various application scenarios increases, the computational cost of high-performing models becomes a significant bottleneck for practical deployment. This has created a pressing need for streaming models that are simultaneously fast (low-latency), lightweight (low computational overhead), and high-fidelity (high audio quality and speaker similarity).
@@
Recent streaming zero-shot VC studies have explored two main architectural paradigms: autoregressive (AR) and non-autoregressive (NAR) frameworks. A representative AR-based streaming VC system is StreamVoice [DBLP:conf/acl/WangCW0W24], which employs a fully causal, context-aware language model. It alternates between semantic bottleneck features and acoustic codec tokens at each AR step, supported by an auxiliary acoustic predictor, to enable streaming zero-shot VC without future look-ahead. Building on this design, StreamVoice+ [DBLP:journals/spl/WangCWXW24] introduces a learnable semantic encoder and a connector, while adapting the backbone via LoRA [DBLP:conf/iclr/HuSWALWWC22]. It further improves quality through a residual-bottleneck connector, forming an end-to-end streaming pipeline. Despite their high-fidelity output, these models fundamentally compromise on the other two critical requirements. Their sequential decoding process inherently leads to high latency, failing the ’fast’ criterion. Moreover, their massive model sizes (101M for StreamVoice and 153M for StreamVoice+) make them far from ’lightweight,’ resulting in prohibitive computational costs that render real-time inference on standard CPUs impractical.
@@
To mitigate the inherent latency of AR models, NAR frameworks have been explored for their potential in parallel generation. DualVC2 [DBLP:conf/icassp/NingJ0WY0B24], for instance, presents a lightweight NAR model based on the conformer architecture, which accelerates inference by incorporating dynamic chunk training. While it achieves faster speeds, its reliance on conventional architectures limits its generalization capability to unseen speakers in zero-shot scenarios. Another NAR approach adapts powerful diffusion models for zero-shot VC. Seed-VC [DBLP:journals/corr/seedvc] employs a diffusion transformer (DiT) [DBLP:conf/iccv/PeeblesX23] architecture within a sliding-window pipeline. Although this mechanism enables streaming VC, the fixed-size windows can disrupt long-term contextual dependencies and speaker characteristics. Thus, while NAR frameworks successfully address the ’fast’ requirement, they often do so by sacrificing ’high-fidelity,’ either through poor generalization in simpler models or context fragmentation in more complex ones. This leaves a clear gap in the field: a framework that can unite the speed of NAR models with the quality and consistency of their AR counterparts.
@@
To address these challenges, we propose MeanVC, a lightweight framework for streaming zero-shot VC that leverages the advantages of both AR and NAR paradigms while addressing their respective limitations. Specifically, MeanVC employs a chunk-wise autoregressive denoising strategy, which processes audio in individual chunks while explicitly conditioning each chunk on the previous one, thereby preserving long-term speaker consistency. To overcome the latency of iterative diffusion samplers, we introduce mean flows [DBLP:journals/corr/MeanFlows], a powerful method that enables the diffusion model to generate high-quality results with 1-NFE (number of function evaluations). We integrate mean flows into the DiT-based decoder. By training the model to regress the average velocity field of the probability flow, MeanVC can synthesize a high-quality spectrogram using 1-NFE. Finally, we incorporate a diffusion adversarial post-training (DAPT) stage to mitigate over-smoothing and further enhance speech quality. With merely 14M parameters, MeanVC achieves superior audio quality and efficiency. Both subjective and objective evaluations demonstrate that MeanVC significantly outperforms existing streaming VC systems. Audio demos and code are publicly available at https://aslp-lab.github.io/MeanVC.
@@ 2.1 Overview
As illustrated in Fig. 1, MeanVC is built on the widely used recognition-synthesis framework [DBLP:journals/taslp/ZhangLD20], consisting of a streaming automatic speech recognition (ASR) module, a speaker encoder, a timbre encoder, a DiT decoder, and a vocoder. First, the pre-trained streaming ASR model extracts bottleneck features (BNFs) from the source waveform. Subsequently, these BNFs are fed into the timbre encoder, which fuses them with fine-grained timbre information extracted from a reference mel-spectrogram to produce timbre BNFs. Meanwhile, speaker embeddings are extracted from a reference waveform via a pre-trained speaker encoder. The DiT decoder generates the converted mel-spectrogram conditioned on both the speaker embeddings and the timbre BNFs. During generation, the source mel-spectrogram is concatenated at the front to serve as a cache, enabling chunk-wise autoregressive denoising and producing the converted spectrogram with the target speaker’s timbre.
@@ 2.2 Chunk-wise Autoregressive Denoising
To enable streaming conversion, we adopt a chunk-wise autoregressive approach that divides the BNFs into smaller chunks and applies a chunk-wise causal mask inspired by MoonCast [DBLP:journals/corr/MoonCast]. This mask strategy allows each chunk to access the historical context from previously generated chunks, ensuring consistency across chunks. Furthermore, we build on conditional flow matching (CFM) to learn optimal transport paths between the two distributions. Given data $x\sim p_{\text{data}}(x)$, prior $\epsilon\sim p_{\text{prior}}(\epsilon)$, the optimal transport flow path can be constructed as: $z_{t}=(1-t)x+t\epsilon$, and the conditional velocity is thus given by $v_{t}=dz_{t}/dt=\epsilon-x$. During training, the objective is to learn a neural network $f_{\theta}$ that minimizes the following CFM loss: $\mathcal{L}_{\text{CFM}}(\theta)=\mathbb{E}_{t,x,\epsilon}\left\|f_{\theta}\left(t,z_{t}\right)-v_{t}\right\|^{2}$.
@@ 2.3 Mean Flows for Generative Modeling
CFM typically requires multiple sampling steps to solve the ordinary differential equation (ODE), and the number of function evaluations directly impacts inference efficiency. With just 1-NFE, the quality of generated spectrograms degrades catastrophically. To address this limitation, we adopt the approach proposed in mean flows [DBLP:journals/corr/MeanFlows], which regresses the average velocity field during training to enable high-quality synthesis with only 1-NFE. Specifically, given a time interval $[r,t]$, the displacement on the ODE trajectory within it is $\int_{r}^{t}v(z_{\tau},\tau)d\tau$. The average velocity is then defined as: $u(z_{t},r,t)\triangleq\frac{1}{t-r}\int_{r}^{t}v(z_{\tau},\tau)d\tau$. By differentiating both sides with respect to $t$ and subsequently rearranging the resulting terms, the relationship between $u$ and $t$ can be derived, yielding the mean flows identity:
@@
By utilizing the Jacobian-vector product (JVP), the total derivative can be expressed as: $\frac{d}{dt}u(z_{t},r,t)=v(z_{t},t)\partial_{z}u+\partial_{t}u$. The velocity $v(z_{t},t)$ is the marginal velocity in flow matching. We follow [DBLP:conf/iclr/LipmanCBNL23] to replace it with the conditional velocity $v_{t}=\epsilon-x$. With this, the target field for training is defined as $u_{\text{tgt}}=v_{t}-(t-r)(v_{t}\partial_{z}u_{\theta}+\partial_{t}u_{\theta})$. We then encourage $f_{\theta}$ to satisfy this field by minimizing the mean flows objective:
@@ 2.4 Diffusion Adversarial Post-Training
Mel-spectrograms generated by flow matching models often suffer from over-smoothing artifacts. While conventional multi-length discriminators [DBLP:journals/corr/HiFiSinger] are employed, we find that high-frequency components in generated mel-spectrograms still exhibit artifacts such as bright lines. Inspired by [DBLP:journals/corr/APT], we leverage the generator architecture as the discriminator backbone to address this issue. Specifically, we employ the DiT model as generator $G$, which generates spectrograms with 1-NFE. The DiT weights are used to initialize the backbone of the discriminator $D$. After training, the discriminator outputs a logit that distinguishes between real samples $x$ and generated samples $\hat{x}$. To achieve this, we modify the DiT architecture by introducing cross-attention-only Transformer blocks at the second and fourth layers of the 4-layer DiT backbone. These blocks use a learnable vector as the query and perform cross-attention with the backbone’s latent representations as keys and values. The outputs are concatenated channel-wise to form a global feature vector, which is then projected to produce a scalar logit output.
@@ 3.1 Experiment Setup
Dataset. We train MeanVC on the open-source Emilia [DBLP:conf/slt/Emilia] corpus. Specifically, we use DNSMOS [DBLP:conf/icassp/Dnsmos] to filter out audio with DNSMOS scores below 3.4, retaining 10,000 hours of mandarin data. All audio files are resampled to 16 kHz for VC training. For zero-shot evaluation, we use the Mandarin subset of the Seed-TTS test set [DBLP:journals/corr/Seed-tts], comprising 2,018 source–target testing pairs. For known speakers evaluation, we fine-tune MeanVC on Aishell3 [DBLP:conf/interspeech/AISHELL-3], selecting four speakers as targets for evaluation. We use 100 clean and 100 noisy clips as source recordings. To extract semantic features, we incorporate the streaming ASR model Fast-U2++ [DBLP:conf/icassp/Fast-U2++], implemented by WeNet [DBLP:conf/interspeech/WeNet] and trained on WenetSpeech [DBLP:conf/icassp/WENETSPEECH].
Implementation Details. MeanVC consists of 14M parameters. The DiT decoder employs four DiT blocks, each with a hidden size of 512 and two attention heads. The timbre encoder incorporates two cross-attention modules, with a hidden size of 256 and four attention heads. Fast-U2++ utilizes a 160 ms chunk size for streaming inference and compresses 16 kHz waveforms into semantic features with a 40 ms frame length. Speaker embeddings are extracted using ECAPA-TDNN [DBLP:conf/interspeech/ECAPA-TDNN]. We employ Vocos [DBLP:conf/iclr/Vocos] as the vocoder to convert mel-spectrograms to high-fidelity 16 kHz speech waveforms.
@@
Baseline Systems. To validate the performance of MeanVC, we compare MeanVC against the following baseline systems: StreamVoice [DBLP:conf/acl/WangCW0W24], a language model-based system that employs fully causal context-awareness. It alternately processes semantic and acoustic features to achieve streaming zero-shot voice conversion; Seed-VC [DBLP:journals/corr/seedvc], a NAR system based on diffusion models. It employs an external timbre shifter to mitigate timbre leakage and achieves streaming zero-shot voice conversion through a sliding-window pipeline; DualVC2 [DBLP:conf/icassp/NingJ0WY0B24], a lightweight NAR system based on the conformer architecture. It uses classic non-causal convolutions and dynamic masking convolutions to achieve streaming any-to-many voice conversion.